# LEMON-Mapping: Loop-Enhanced Large-Scale Multi-Session Point Cloud Merging and Optimization for Globally Consistent Mapping

### v1

@@ III-B Loop Processing Module
We implement a region-growing algorithm based on Breadth-First Search (BFS). Starting from the spatial center of each loop, the algorithm incrementally searches for nearby loops within a predefined radius. If neighboring loops are found, the search expands from their centers, continuing recursively until no additional nearby loops are detected. Loops that form such spatial clusters are labeled as clustered loops, while those without adjacent loops are marked as isolated loops. An illustrative example of this classification process is shown in Fig. 2.
Loop Recall.
@@
Since the outlier rejection step adopts strict criteria to ensure robustness and accuracy, it may inadvertently discard valid loop closures as, resulting in false negatives. These missing constraints between robot trajectories hinder the correction of divergence in overlapping areas, potentially degrading the consistency of the fused multi-robot map.
@@
To address this issue, we propose a loop recall mechanism to recover previously discarded valid loops. After the first pose graph optimization (Section III-D) aligns all robot trajectories into the same coordinate system using correct loops, the updated poses are used to re-evaluate previously rejected loops. If the Euclidean distance between the associated poses is below a threshold (2 meters in our system), the loop is recalled for further optimization. This lightweight and distance-based strategy effectively recovers useful constraints and enhances subsequent optimization.
III-C Spatial Bundle Adjustment
@@
Traditional bundle adjustment methods [23, 24, 25] jointly optimize sequential poses within a sliding window of temporally ordered data. However, they struggle with scenarios involving revisiting the same location over long time spans, making them unsuitable for multi-robot systems. Different from previous BA methods, our spatial bundle adjustment operates within spatially localized windows, jointly optimizing poses from different robots in overlapping regions simultaneously. This design effectively reduces map divergence across sessions, as demonstrated in Fig.3 and Fig.8. Our approach specifically focuses on loop closure regions for two reasons: (1) these regions often exhibit significant spatial overlap among submaps and suffer from high divergence; (2) they contain multiple observations of the same geometric structure from different times or agents, offering rich constraints for accurate joint optimization.
@@
The spatial bundle adjustment is performed on all correct loops after the first pose graph optimization and the loop recall step. All available loops are divided into two categories based on the principle of loop classification in Section III-B. The poses of all robots after FPGO are built into a kd-tree using their spatial positions for faster searching. For each loop, we define a spherical spatial window centered at the midpoint between the two poses involved in the loop. Within this window, all nearby poses are included in the spatial bundle adjustment. For faster searching, we employ a radius-based search using the pose kd-tree to efficiently retrieve the poses in the spherical region. These selected poses form the local optimization window for the spatial bundle adjustment, ensuring that only the spatially correlated data around the loop are refined. Depending on the type of loop closure classified in Section III-B, we propose two forms of spatial bundle adjustment designed for their respective characteristics. For isolated loops, we apply our developed spatial diffusion bundle adjustment (DBA). For clustered loops, a spatial variant of the hierarchical bundle adjustment (HBA) [25] is proposed.
@@
DBA for isolated loops.
Our DBA is built upon BALM2 [24], which jointly optimizes poses by minimizing the distances from feature points to the corresponding planes. However, BALM, which relies on accurate initial plane estimation, struggles or fails to converge reliably when the same plane observed by different robots appears misaligned due to map divergence. In such cases, the same structure may be voxelized as separate multi-layer planes (shown in Fig. 3 (a)), making direct optimization invalid. To address this, we partition the poses within the spatial window by robot into distinct clusters, each with its own point cloud, as illustrated in the left part of Fig.3. We then apply GICP [41] to roughly align these clusters, reducing plane stratification and enabling reliable plane estimation for subsequent optimization.
After the initial pose graph optimization incorporating an isolated loop, the poses near the loop center are effectively aligned and optimized. However, as the distance from the loop center increases, the influence of the loop constraints in FPGO gradually weakens, leading to less accurate optimization and alignment of distant poses. Therefore, for isolated loops, the key point is to figure out how large its influence range is and align the surrounding affected poses. We develop spatial Diffusion Bundle Adjustment (DBA) to tackle this issue.
In DBA, we begin with a set of LiDAR poses confined to a small range near an isolated loop. We then gradually expand the set by diffusely incorporating additional LiDAR poses from increasingly wider ranges. The key principle is that, within the loop influence region, poses located further from the loop center contribute less accurately to the representation of the loop’s features. The details of DBA are as follows:
@@
where it does not change when optimizing ${}^{1}\mathbf{T}$ compared with jointly solving the BA. The same is true for the second-order gradient, so we only need to freeze the update for ${}^{0}\mathbf{T}$ poses (${{}^{0}\Delta\mathbf{T}}=0$). The solver that jointly solves this BA problem has a complexity of $O\left(M_{f}M_{p}+M_{f}M_{p}^{2}+M_{p}^{3}\right)$, while by masking the Jacobian and Hessian flexibly, our DBA could achieve a complexity of:
@@
HBA for clustered loops.
Figure 4: (a) shows the Multi-robot poses in a spatial window. (b) shows the reordered poses using PCA for HBA.
For two trajectories from different robots with significant overlapping regions, loop closures tend to appear densely, and adjacent loops often share large overlapping areas in their associated point cloud maps. In such clustered loop groups, it is crucial to jointly process the loops within the entire overlapping region to maintain local consistency and optimization accuracy. If we optimize the poses by processing loop closures sequentially and independently rather than jointly, this may introduce inconsistencies and even divergence in overlapping regions due to contradictory constraints. To address this, we adapt and extend HBA [25] into a multi-robot framework for clustered loop scenarios.
The original HBA relies on temporally ordered poses from a single robot, assuming adjacent poses share common point cloud features. However, in our multi-robot setting, poses selected via kd-tree radius search within a loop cluster are unordered and from different robots. To address this, we use Principal Component Analysis (PCA) to analyze the spatial distribution of selected poses and reorder them along the principal axis defined by the largest eigenvalue, as illustrated in Fig. 4. This spatial reordering ensures neighboring poses in the optimization sequence are also spatially close, promoting shared features and effective constraints. Once reordered, all poses are processed consistently in a unified optimization window, enabling the enhanced HBA to accurately align the geometry across overlapping regions from multiple robots.
III-D Two-Step Pose Graph Optimization
The two-step pose graph optimization are used to align and refine the global structure of multi-robot trajectories. The details of each step are as follows.
@@ III-D Two-Step Pose Graph Optimization
First Pose graph optimization (FPGO).
@@
Following loop processing, we perform the first pose graph optimization to coarsely align the trajectories of all robots. We construct a centralized pose graph shown in Fig. 5, in which each robot’s pose sequence is included as a subgraph. The graph incorporates three types of constraint: odometry, self-loop, and inter-loop. The first pose of the first robot is fixed to define the origin point of the world coordinate, ensuring that all other robot trajectories are aligned to this global reference frame. With this initial alignment established, we acquire a rough estimate of the relative positions among all robot trajectories, which enables us to recover the valid but previously rejected loops (in Section III-B).
@@
Figure 6: The upper part of the picture shows the sparsification of BA constraints in a spatial window at loop. The sparse BA constraints are add to pose graph below along with odometry constrains.
Last Pose Graph Optimization (LPGO).
@@
Although spatial BA (Section III-C) improves the accuracy of local pose near loop closures, it does not propagate this accuracy globally and may break the continuity of odometry, resulting in even worse global consistency. We develop the last pose graph optimization to address this, and the details are as follows.
@@
To achieve global consistency while preserving local accuracy, we design the final pose graph optimization with two types of constraints: odometry constraints to maintain trajectory smoothness, sparsified BA-based constraints to preserve refined local structures (see Fig. 6). For each robot, odometry constraints are added between adjacent poses to preserve temporal continuity. In cases where a BA-optimized pose is adjacent to an under-optimized one, a high-weight odometry constraint is applied to mitigate potential discontinuity introduced by local optimization. Conversely, lower weights are assigned between under-optimized poses to allow greater global adjustment flexibility. Additionally, pose pairs within the spatial BA window that exhibit strong geometric overlap and reliable correspondence are assigned high-weight BA constraints to preserve local accuracy. The specific selection criteria for these constraints are detailed in the following paragraphs.
@@
We use two different metrics for constraint selection between pose pairs of different robots and pose pairs of the same robot. For pose pairs across different robots, we evaluate the similarity of their descriptors obtained from RING++ [21] and add BA constraints only when the similarity exceeds a specified threshold. For pose pairs within the same robot, we apply pairwise registration techniques from [43] to selectively retain constraints between poses that significantly restrain each other. The sparsification process using pairwise registration is as follows.
For an edge $\mathbf{E}_{k}$ between two pose nodes, the $\mathbf{E}_{k}^{0}$-th pose and the $\mathbf{E}_{k}^{1}$-th pose, of the same robot. The relative pose residual ($\bm{\epsilon}_{k}$) and covariance ($\mathbf{\Omega}_{k}$) between these poses can be approximated using the pairwise registration formulation. The $\mathbf{E}_{k}^{0}$-th pose and the $\mathbf{E}_{k}^{1}$-th pose are denoted as $\mathbf{R}_{L_{0}},\mathbf{t}_{L_{0}}$ and $\mathbf{R}_{L_{1}},\mathbf{t}_{L_{1}}$ respectively for convenience. The relative pose is denoted as $\mathbf{R}^{L_{0}}_{L_{1}}=\mathbf{R}^{\top}_{L_{0}}\mathbf{R}_{L_{1}},\mathbf{t}^{L_{0}}_{L_{1}}=\mathbf{R}^{\top}_{L_{0}}(-\mathbf{t}_{L_{0}}+\mathbf{t}_{L_{1}})$. Utilizing nearest searching and taking the point-to-point correspondences $\{(\mathbf{P}^{L_{0}}_{u},\mathbf{P}^{L_{1}}_{u})\}^{U}_{u=1}$ as input, the registration function $\bm{\epsilon}^{\mathrm{reg}}_{k}$ for $\mathbf{E}_{k}$ is as follows.
@@
The Jacobian of $\bm{\epsilon}^{\mathrm{reg}}_{k}$ respect to the relative pose $\mathbf{R}^{L_{0}}_{L_{1}},\mathbf{t}^{L_{0}}_{L_{1}}$ is calculated by (15).
@@
The covariance $\mathbf{\Omega}_{k}$ of the registration function $\bm{\epsilon}^{\mathrm{reg}}_{k}$ is calculated as follows.
@@
IV Experiments
IV-A Experimental Setup
All algorithms are implemented in C++ using the Robot Operating System (ROS). To evaluate the performance of the proposed multi-session map merging framework, we conduct experiments on several datasets, including the publicly available S3E [44], GEODE [45], MARS-LVIG [1], R3LIVE [46] dataset, as well as a self-collected dataset. The GEODE, MARS-LVIG and R3LIVE datasets are segmented into multiple sessions with overlapping areas, but the starting point of each part is different and the relative transformation is unknown. The parameters of the public datasets and our self-collected dataset are shown in Table II and Table I, respectively. For all public datasets, initial odometry is obtained using the LiDAR-Inertial Odometry method recommended by the respective dataset, while the initial odometry for our own dataset is generated using FAST-LIO2 [14].
We first evaluate our proposed spatial bundle adjustment by comparing it against the conventional LiDAR-based bundle adjustment methods BALM2 [24] and HBA [25] (Section IV-B). Next, we assess the performance of our complete multi-session merging framework against DCL-SLAM [27] and LAMM [29] (Section IV-C). An ablation study is then conducted to analyze the contributions of individual components within our framework (Section IV-D). Finally, we test the scalability of our method in handling a large number of sessions (Section IV-E).
@@ IV-A Experimental Setup
Dataset Environment Trajectory Length (m) LiDAR
@@
Yard Iutdoor & Outdoor 232 Mid360
Laboratory Indoor & Outdoor 98 Mid360
TABLE II: OVERVIEW OF OPEN DATASET
Dataset Sequence Robot Num Traj Length (m) LiDAR
@@
LVIG Airport 2 4135 Avia
@@
HKU Mainbuilding 20 945 Avia
Figure 7: Result of single robot study. (a) shows the global map generated by our method. (b)-(d) shows the details of the local map at loop.
IV-B Single-Robot Study
Each sequence in our dataset includes loop closures that enable spatial BA (Section III-C). We compare our spatial BA with BALM2 [24] and HBA [25], using average $z$-axis drift ($z$-DRIFT) and $z$-axis Root Mean Square Error ($z$-RMSE) as primary evaluation metrics. Both metrics are computed with respect to the $z$-value of the first frame as a reference, allowing a consistent evaluation of vertical alignment over time. Due to the absence of ground truth, we also compute the Mean Map Entropy (MME) using MapEval [47], where lower values indicate better map consistency and reduced clutter. To ensure a fair comparison of the performance of the three BA methods, all methods are evaluated using raw odometry trajectories without any prior loop-based refinement. For computational feasibility in large-scale scenes, BALM2 is executed in a sliding-window configuration.
TABLE III: MME, $z$-DRIFT AND $z$-RMSE OF SINGLE ROBOT STUDY
       Scene Method MME $z$-Drift $z$-RMSE
@@ IV-B Single-Robot Study
IV-C Multi-Robot Study
TABLE IV: RMSE OF THE ATE(m) OF OPEN DATASET
         Scene DCL-SLAM LAMM Ours

### v2

@@ IV-B Loop Classification
We implement a region-growing algorithm based on Breadth-First Search (BFS). Starting from the spatial center of each loop, the algorithm incrementally searches for nearby loops within a predefined radius. If neighboring loops are found, the search expands from their centers, continuing recursively until no additional nearby loops are detected. Loops that form such spatial clusters are labeled as clustered loops, while those without adjacent loops are marked as isolated loops. An illustrative example of this classification process in the S3E Campus_1 dataset is shown in Fig. 2.
IV-C Loop Recall
@@
Since the outlier rejection step adopts strict criteria to ensure robustness and accuracy, it may inadvertently discard valid loop closures, thereby resulting in false negatives. These missing constraints between robot trajectories hinder the correction of divergence in overlapping areas, potentially degrading the consistency of the fused multi-robot map.
@@
To address this issue, we propose a loop recall mechanism to recover previously discarded valid loops. After the first pose graph optimization (Section VI) aligns all robot trajectories into the same coordinate system using valid loops, the updated poses are used to re-evaluate previously rejected loops. If the Euclidean distance between the associated poses is below a threshold (2 meters in our system), the loop is recalled for further optimization. This lightweight and distance-based strategy effectively recovers useful constraints and enhances subsequent optimization. Fig. 12 demonstrates the impact of loop recall on the merging of point cloud maps.
V Spatial Bundle Adjustment
@@
Traditional BA methods [23, 24, 25] jointly optimize sequential poses within a sliding window of temporally ordered data. However, they struggle with scenarios involving revisiting the same location over long time spans, making them unsuitable for multi-robot systems. Different from previous BA methods, our spatial BA jointly optimizes poses from different robots in local spatial windows simultaneously. This design effectively reduces map divergence across sessions, as demonstrated in Fig.3 and Fig.8. Our approach specifically focuses on loop closure regions for two reasons: (1) these regions often exhibit significant spatial overlap among submaps and suffer from high divergence; (2) they contain multiple observations of the same geometric structure from different times or agents, offering rich constraints for accurate joint optimization.
@@
The spatial BA is performed on all correct loops after the FPGO and the loop recall step. All available loops are divided into two categories based on the principle of loop classification in Section IV. The poses of all robots are built into a kd-tree using their spatial positions for faster searching. For each loop, we define a spherical spatial window centered at the midpoint between the two poses involved in the loop closure. We employ a radius-based search using the pose kd-tree to efficiently retrieve the poses in the spherical region. These selected poses form the local optimization window for the spatial BA, ensuring that only the spatially correlated data around the loop are refined. Depending on the type of loop closure classified in Section IV, we propose two forms of spatial bundle adjustment designed for their respective characteristics to refine multi-robot maps. For isolated loops, we apply our developed spatial diffusion bundle adjustment (DBA). For clustered loops, a spatial variant of hierarchical bundle adjustment (HBA) is proposed.
@@
V-A DBA for Isolated Loops
Our DBA is built upon BALM2 [24]. However, it relies on accurate initial plane estimation, which may fail to converge reliably when the same plane observed by different robots appears misaligned due to map divergence. In such cases, the same structure may be voxelized as separate multi-layer planes (shown in Fig. 3 (a)), making direct optimization invalid. To address this, we partition the poses within the spatial window by robot into distinct clusters, each with its own point cloud, as illustrated in the left part of Fig.3. We then apply GICP [42] to roughly align these clusters, reducing plane stratification and enabling reliable plane estimation for subsequent optimization. After the initial pose graph optimization incorporating an isolated loop, the poses near the loop center are effectively aligned and optimized. However, as the distance from the loop center increases, the influence of loop constraints in FPGO gradually weakens, leading to less accurate optimization and alignment of distant poses. Therefore, for an isolated loop, the key point is to figure out how large its influence range is and align the surrounding affected poses. We develop DBA to tackle this issue.
We begin with a set of LiDAR poses confined to a small range near an isolated loop. We then gradually expand the set by diffusely incorporating additional LiDAR poses from increasingly wider ranges. The key principle is that, within the loop influence region, poses located further from the loop center contribute less accurately to the representation of the loop’s features. The details of DBA are as follows.
@@ V-A DBA for Isolated Loops
where it does not change when optimizing ${}^{1}\mathbf{T}$ compared with jointly solving the BA. The same is true for the second-order gradient. Based on (5), we only need to freeze the update for ${}^{0}\mathbf{T}$ poses (${{}^{0}\Delta\mathbf{T}}=0$).
The solver that jointly solves this BA problem has a complexity of $O\left(M_{f}M_{p}+M_{f}M_{p}^{2}+M_{p}^{3}\right)$, while by masking the Jacobian and Hessian flexibly, our DBA could achieve a complexity of:
@@
V-B HBA for Clustered Loops
Figure 4: (a) shows the multi-robot poses in a spatial window. (b) shows the reordered poses using PCA for spatial HBA.
For two trajectories from different robots with significant spatial overlap, loop closures tend to appear densely, and the associated point cloud maps of adjacent loop closures often share large common areas, thus forming a clustered loop groups. In such clustered loop groups, it is crucial to jointly process the loops within the entire overlapping region to maintain local consistency. If we optimize the poses by processing loop closures sequentially and independently, this may introduce inconsistency or even divergence in overlapping regions due to contradictory constraints. To address this, we adapt and extend HBA into a multi-robot framework for clustered loop scenarios.
The original HBA relies on temporally ordered poses from a single robot, assuming that adjacent poses share common point cloud features. However, in our multi-robot setting, poses selected via kd-tree radius search within a loop cluster are unordered and from different robots. To address this, we use Principal Component Analysis (PCA) [44] to analyze the spatial distribution of selected poses and reorder them along the principal axis defined by the largest eigenvalue, as illustrated in Fig. 4. This spatial reordering ensures that neighboring poses in the optimization sequence are also spatially close, promoting shared features and effective constraints. Once reordered, all poses are processed consistently in a unified optimization window, allowing the enhanced HBA to accurately align the geometry across overlapping regions from multiple robots.
VI Two-Step Pose Graph Optimization
The two-step pose graph optimization is used to align and refine the global structure of multi-robot trajectories. The details of each step are as follows.
@@ VI Two-Step Pose Graph Optimization
VI-A First Pose Graph Optimization
@@
Following loop processing, we perform the first pose graph optimization (FPGO) to coarsely align the trajectories of all robots. We construct a centralized pose graph shown in Fig. 5, in which each robot’s pose sequence is included as a subgraph. The graph incorporates three types of constraint: odometry, self-loops, and inter-loops. The first pose of the first robot is fixed to define the origin point of the world coordinate, ensuring that all other robot trajectories are aligned to this global reference frame. With this initial alignment established, we acquire a rough estimate of the relative positions among all robot trajectories, which enables us to recover the valid but previously rejected loops (in Section IV).
@@
Figure 6: The upper part of the picture shows the sparsification of BA constraints in a spatial window at loop. The sparse BA constraints are add to pose graph below along with odometry constrains with different weights.
VI-B Last Pose Graph Optimization
@@
Although spatial BA (Section V) improves the accuracy of local pose near loop closures, it does not propagate this accuracy globally and may break the continuity of odometry, resulting in even worse global consistency. To address this issue, we develop the last pose graph optimization (LPGO), and the details are as follows.
@@
We design the LPGO with two types of constraints: odometry constraints to maintain trajectory smoothness, and sparsified BA-based constraints to preserve refined local structures (see Fig. 6). For each robot, odometry constraints are added between adjacent poses to preserve temporal continuity. In cases where a BA-optimized pose is adjacent to an unoptimized one, a high-weight odometry constraint is applied to mitigate potential discontinuities introduced by local optimization. In contrast, lower weights are assigned between two unoptimized poses to allow greater global adjustment flexibility. Additionally, pose pairs within the spatial BA window that exhibit strong geometric overlap and reliable correspondence are assigned high-weight BA constraints to preserve local accuracy. The specific selection criteria for these constraints are detailed in the following paragraphs.
@@
We use two different metrics for constraint selection between pose pairs of different robots and pose pairs of the same robot. For pose pairs across different robots, we evaluate the similarity of their descriptors obtained from RING++ [21] and add BA constraints only when the similarity exceeds a specified threshold. For pose pairs within the same robot, we apply pair-wise registration techniques from [45] to selectively retain constraints between poses that significantly constrain each other. The sparsification process using pair-wise registration is as follows.
For an edge $\mathbf{E}_{k}$ connecting two pose nodes belonging to the same robot, let $\mathbf{E}_{k}^{0}$ and $\mathbf{E}_{k}^{1}$ denote the indices of the corresponding poses. The residual $\bm{\epsilon}_{k}$ and its associated covariance matrix $\mathbf{\Omega}_{k}$ between the two poses can be estimated using a pair-wise registration approach. For clarity, we denote the poses at $\mathbf{E}_{k}^{0}$ and $\mathbf{E}_{k}^{1}$ as $(\mathbf{R}_{L_{0}},\mathbf{t}_{L_{0}})$ and $(\mathbf{R}_{L_{1}},\mathbf{t}_{L_{1}})$, respectively. The relative transformation from frame $L_{0}$ to $L_{1}$ is computed as $\mathbf{R}^{L_{0}}_{L_{1}}=\mathbf{R}_{L_{0}}^{\top}\mathbf{R}_{L_{1}}$ and $\mathbf{t}^{L_{0}}_{L_{1}}=\mathbf{R}_{L_{0}}^{\top}(\mathbf{t}_{L_{1}}-\mathbf{t}_{L_{0}})$. Based on nearest neighbor matching, we obtain point-to-point correspondences $\{(\mathbf{P}^{L_{0}}_{u},\mathbf{P}^{L_{1}}_{u})\}_{u=1}^{U}$, where $U$ denotes the total number of matched pairs. These correspondences are used to formulate the registration residual function $\bm{\epsilon}^{\mathrm{reg}}_{k}$ associated with edge $\mathbf{E}_{k}$.
@@
The Jacobian of $\bm{\epsilon}^{\mathrm{reg}}_{k}$ respect to the two relative poses of the same robot $\mathbf{R}^{L_{0}}_{L_{1}},\mathbf{t}^{L_{0}}_{L_{1}}$ is calculated by (15).
@@
The covariance $\mathbf{\Omega}_{k}$ of the registration function $\bm{\epsilon}^{\mathrm{reg}}_{k}$ can be calculated as follows.
@@
VII Experiments
VII-A Experimental Setup
All algorithms are implemented in C++ using the Robot Operating System (ROS) [46]. To evaluate the performance of the proposed multi-session map merging framework, we conduct experiments on several datasets, including the publicly available S3E [41], GEODE [47], MARS-LVIG [1], R3LIVE [48] dataset, as well as a self-collected dataset. The GEODE, MARS-LVIG and R3LIVE datasets are segmented into multiple sessions with overlapping areas, but the starting point of each part is different and the relative transformation is unknown. The parameters of our self-collected dataset and the public datasets are shown in Table I and Table II, respectively. For all public datasets, initial odometry is obtained using the LiDAR-Inertial-Odometry method recommended by the respective dataset, while the initial odometry for our own dataset is generated using FAST-LIO2 [14].
We first evaluate our proposed spatial bundle adjustment by comparing it with the conventional LiDAR-based bundle adjustment methods BALM2 [24] and HBA [25] (Section VII-B). Next, we assess the performance of our complete multi-session merging framework against DCL-SLAM [27] and LAMM [29] (Section VII-C). An ablation study is then conducted to analyze the contributions of individual components within our framework (Section VII-D). Finally, we test the scalability of our method in handling a large number of sessions (Section VII-E).
@@ VII-A Experimental Setup
Sequence Environment Trajectory Length (m) LiDAR
@@
Yard Indoor & Outdoor 232 Mid360
Laboratory Outdoor 98 Mid360
Flying Arena Indoor 280 Mid360
TABLE II: OVERVIEW OF PUBLIC DATASET
Dataset Sequence Robot Number Trajectory Length (m) LiDAR
@@
Offroad 2 542 Ouster
MARS-LVIG Airport 2 4135 Avia
@@
HKUST Campus 20 945 Avia
Figure 7: Result of single-robot study in our Garage dataset. (a) shows the global map generated by our method. (b)-(d) show the details of the local maps near the loop closure for different methods.
TABLE III: MME, $z$-DRIFT AND $z$-RMSE OF SINGLE-ROBOT STUDY
       Sequence Method MME $z$-Drift $z$-RMSE
@@
Figure 8: The map merging result of S3E library. The local maps generated by three methods are selected for comparison. Our method almost eliminates the divergence of local multi-session point cloud maps. While LAMM [29] and DCL-SLAM [27] suffer from serious inconsistencies.
VII-B Single-Robot Study
Each sequence in our dataset includes loop closures that enable spatial BA (Section V). We compare our spatial BA with BALM2 [24] and HBA [25], using average $z$-axis drift ($z$-DRIFT) and $z$-axis Root Mean Square Error ($z$-RMSE) as primary evaluation metrics. Both metrics are computed with respect to the $z$-value of the first frame as a reference, allowing a consistent evaluation of vertical alignment over time. Due to the absence of ground truth, we also compute the Mean Map Entropy (MME) using MapEval [49], where lower values indicate better map consistency and reduced clutter. To ensure a fair comparison of the performance of the three BA methods, all methods are evaluated using raw odometry trajectories without any prior loop-based refinement. For computational feasibility in large-scale scenes, BALM2 is executed in a sliding-window configuration.
Fig. 7 shows the mapping performance of HBA, BALM2 and our proposed method in the Garage sequence, with the framed area indicating the region where loop closures occur. It can be seen that both HBA and BALM2 exhibit significant layering in this area, whereas our method achieves superior alignment. This result clearly demonstrates the strong capability of our spatial BA approach in handling loop closure regions.
Figure 9: The optimized trajectory, position, and position error for proposed method in MARS-LVIG Island ((a)-(c)) and Town ((d)-(f)).
Figure 10: Comparison of multi-robot trajectories and error between our method, LAMM, and DCL-SLAM on S3E Library. The ground truth is shown as a dashed line in the figure.
Figure 11: The same comparison of multi-robot trajectories on S3E Tunnel as Fig. 10.
VII-C Multi-Robot Study
TABLE IV: RMSE OF THE ATE(m) OF MULTI-ROBOT STUDY
         Sequence DCL-SLAM LAMM Ours